# Distributed Online Feedback Optimization for Real-time Distribution System Voltage Regulation

### v2

@@ Abstract
We investigate the real-time distribution system voltage regulation problem employing online feedback optimization (OFO) and short-range communication between physical neighbours. OFO does not need an accurate grid topology nor estimated consumption of non-controllable loads, affords fast calculations, and demonstrates robustness to uncertainties due to its feedback-based nature. These characteristics render it particularly suitable for real-time distribution system applications. However, many OFO controllers require centralized communication, which lack robustness to single-point failures. This paper proposes a distributed OFO design based on a nested feedback optimization strategy and analyzes its convergence. Numerical results reveal that the proposed design achieves satisfactory voltage regulation and outperforms other distributed approaches.
@@
Distributed communication, online feedback optimization, voltage regulation, reactive power control
@@ I Introduction
With the proliferation of distributed energy resources (DERs) in distribution systems, real-time voltage regulation has become a challenging task for distribution system operators (DSOs). While local control cannot guarantee the desired voltage regulation [1], model-based feedforward optimization requires an accurate grid model, load measurements, and lacks robustness to system uncertainties and disturbances [2].
To overcome these limitations, online feedback optimization (OFO) has been proposed as a promising solution approach. OFO utilizes measurements as feedback and employs optimization algorithms as feedback controllers to steer physical systems towards their optimal operating points [2]; see Fig. 1 for comparison between feedforward and feedback systems. The feedback-based implementation eliminates the need for an accurate grid model and real-time load data and brings robustness to OFO. Compared to deep reinforcement learning, OFO does not need a complicated training process and has guarantees to pursue grid-level objectives. Compared to using traditional voltage regulators such as on-load tap changers and capacitor banks, OFO provides fast voltage regulation and can complement these slow-reacting devices in real time [3].
While many OFO controllers rely on centralized communication [4, 5, 6, 7, 8], distributed OFO based on communication between physical neighbours appears a compelling approach due to its robustness to single-point failures[9, 1, 3, 10, 11, 12]. Among the existing distributed designs, [9, 1, 3] demand particular forms of objective functions, while [10] relies on disseminating global information through neighboring nodes. However, the systematic delays inherent in this information propagation may raise stability issues, potentially undermining its efficacy in practical distribution systems. In [11], the distributed design was achieved by only keeping components in the network sensitivity matrix related to local and neighbouring nodes. This however does not assure voltage regulation.
In this paper, we propose a distributed OFO design related to that in [12], where the distributed implementation was enabled by the sparsity of the inverse of the network sensitivity matrix $\mathbf{X}$, i.e. $\mathbf{X}^{-1}$ is sparse. Both approaches continue with scaling the gradient by $\mathbf{X}^{-1}$ followed by a projection step to ensure satisfaction of the DER capacity constraint. It is conceptually attractive to use the Euclidean distance in the projection, which can be implemented using only local information. However, as shown in [13], the distance measured by the norm $\lVert\mathbf{z}\rVert_{\mathbf{X}}=\sqrt{\mathbf{z}^{\intercal}\mathbf{X}\mathbf{z}}$ should be used instead in the projection step to ensure descent iterations and thus algorithm convergence. With the Euclidean distance, it would be a two-metric approach which can lead to algorithm divergence [13]. Solving this non-Euclidean projection problem would however require global information since $\mathbf{X}$ is dense, contradicting the distributed OFO design. As a remedy, [12] introduced additional dualization of the DER capacity limits on top of the Euclidean projection to ensure the DER constraint satisfaction. Shown in their numerical result and our simulation, this strategy takes a prohibitively large number of iterations to converge, which jeopardizes its performance in online implementation.
@@
Figure 1: Block diagrams of feedforward and feedback systems for voltage regulation using reactive power, where $q$ represents system input (reactive power), $v$ represents system output (voltages), $\omega$ ($\widehat{\omega}$) represents (estimated) system disturbance (non-controllable loads), and $h$ maps system input and disturbance to system output.
@@
We tackle this challenge by proposing an iterative algorithm to solve this non-Euclidean projection problem that requires only local information and operates in an online feedback-based manner. This leads to our overall nested online feedback optimization approach, where the outer loop comprises the OFO iterations requiring communication between physical neighbours and yielding tentative but not necessarily feasible DER setpoints, and the inner loop solves the non-Euclidean projection problem to map the tentative setpoints to actual feasible setpoints using local information.
@@
The remainder of this paper is structured as follows: Section II presents the problem formulation and the proposed nested approach and analyzes its convergence. Section III presents simulation results, while Section IV draws conclusions and discusses future work.
II Online feedback optimization
II-A Problem formulation
@@
Consider the distribution system voltage regulation problem using reactive power as formulated in (1). DERs are located in the set $\mathcal{N}$, and $q_{i}$ represents their respective reactive power injection. The objective function (1a) minimizes the reactive power use cost, where $c_{i}>0$ is the cost parameter. Constraint (1b) enforces the lower and upper voltage limits $\underline{v}$ and $\overline{v}$, where $\mu_{i}$ and $\lambda_{i}$ are the respective dual variables. Finally, (1c) ensures that the reactive power injection $q_{i}$ respects the lower and upper limits $\underline{q}_{i}$ and $\overline{q}_{i}$. Note that $v_{i}$ is dependent on $q_{i}$ and other disturbances in the system such as active and reactive power consumption of non-controllable load units.
$\underset{q_{i},\forall i\in\mathcal{N}}{\operatorname{minimize}}$ $\quad f:=\frac{1}{2}c_{i}q_{i}^{2}$ (1a)
@@
$\operatorname{s.t.}$ $\quad\underline{v}\leq v_{i}\leq\overline{v}:\mu_{i},\lambda_{i},\forall i\in\mathcal{N}$ (1b)
$\underline{q}_{i}\leq q_{i}\leq\overline{q}_{i},\forall i\in\mathcal{N}$ (1c)
II-B Centralized primal-dual gradient projection algorithm
@@
Solving (1) offline in a feedforward way demands an accurate grid model and disturbance data to evaluate $v_{i}$, which might not be available. The feedforward approach also lacks robustness [2]. Thus, this section introduces an instance of OFO based on the primal-dual gradient projection (PDGP) algorithm [4, 5] to track the time-varying optimizer of (1), which requires a centralized gather-and-broadcast communication architecture. The PDGP algorithm performs projected gradient ascent and descent steps iteratively for the dual and primal variables, respectively. At each time step (iteration) $k$, it repeats the following three steps:
@@
    For each bus $i\in\mathcal{N}$, collect its voltage measurement $\tilde{v}_{i}^{k}$ and update $\lambda_{i}$ and $\mu_{i}$, where the projection operator is defined as $[u]^{+}:=\max(u,0)$, $\alpha$ with various superscripts represents step sizes, while $r^{d}$ and $r^{p}$ in (2c) are the dual and primal regularization factors [4, 5], respectively.
    $\lambda_{i}^{k+1}\leftarrow\big{[}\lambda_{i}^{k}+\alpha^{d}(\tilde{v}_{i}^{k}-\overline{v}-r^{d}\lambda_{i}^{k})\big{]}^{+}$ (2a)
    $\mu_{i}^{k+1}\leftarrow\big{[}\mu_{i}^{k}+\alpha^{d}(\underline{v}-\tilde{v}_{i}^{k}-r^{d}\mu_{i}^{k})\big{]}^{+}$ (2b)
@@
    The DSO broadcasts the updated dual variables $\boldsymbol{\lambda}^{k+1}$ and $\boldsymbol{\mu}^{k+1}$ to all buses, where $\boldsymbol{\lambda}^{k+1}$ and $\boldsymbol{\mu}^{k+1}$ are vectors constructed by stacking $\lambda_{i}^{k+1}$ and $\mu_{i}^{k+1}$, respectively.
@@
    For each bus $i\in\mathcal{N}$, update its reactive power setpoints $q_{i}$, where $[u]_{\underline{q}_{i}}^{\overline{q}_{i}}$ represents the Euclidean projection into $[\underline{q}_{i},\overline{q}_{i}]$, $\mathbf{X}$ is a $|\mathcal{N}|$-dimensional dense matrix mapping reactive power injections to voltages, i.e. the sensitivity matrix, as constructed in [12], and $\mathbf{q}^{k}$ is the stack of $q^{k}_{i}$. Since $\mathbf{X}$ is dense, Step 2 is needed to provide information for Step 3.
    $q_{i}^{k+1}\leftarrow\big{[}q_{i}^{k}-\alpha[c_{i}q_{i}^{k}+[\mathbf{X}(\boldsymbol{\lambda}^{k+1}-\boldsymbol{\mu}^{k+1}+r^{p}\mathbf{q}^{k})]_{i}]\big{]}_{\underline{q}_{i}^{k}}^{\overline{q}_{i}^{k}}$ (2c)
II-C Two-metric distributed approach
@@
To avoid the centralized communication requirement and enable the distributed communication architecture where communication is only between physical neighbours, [12] proposed to scale the gradient in (2c) by a sparse positive definite matrix $\mathbf{X}^{-1}$, which has non-zero entries only in the diagonal and the $i$-th row $j$-th column if buses $i$ and $j$ are adjacent. Continuing with the Euclidean projection yields (3) which replaces (2c) in the OFO where $\mathbf{C}$ is a $|\mathcal{N}|$-dimensional diagonal matrix with $c_{i}$ defined in (1a).
$q_{i}^{k+1}\leftarrow\big{[}q_{i}^{k}-\alpha[[\mathbf{X}^{-1}\mathbf{C}\mathbf{q}^{k}]_{i}+\lambda_{i}^{k+1}-\mu_{i}^{k+1}+r^{p}q_{i}^{k}]\big{]}_{\underline{q}_{i}^{k}}^{\overline{q}_{i}^{k}}$ (3)
Computing (3) for node $i$ now requires only information exchange with its physical neighbours. We stack $q_{i}^{k}-\alpha[[\mathbf{X}^{-1}\mathbf{C}\mathbf{q}^{k}]_{i}+\lambda_{i}^{k+1}-\mu_{i}^{k+1}+r^{p}q_{i}^{k}]$ for all nodes (tentative but not necessarily feasible setpoints to be projected) as the vector $\mathbf{\dot{q}}^{k+1}$ and define $\mathcal{X}^{k}$ as $[\underline{q}_{1}^{k},\overline{q}_{1}^{k}]\times[\underline{q}_{2}^{k},\overline{q}_{2}^{k}]\times\cdots\times[\underline{q}_{|\mathcal{N}|}^{k},\overline{q}_{|\mathcal{N}|}^{k}]$, where $\times$ is the Cartesian product of sets.
II-D Proposed distributed approach

### v3

@@ Abstract
We investigate the real-time voltage regulation problem in distribution systems employing online feedback optimization (OFO) with short-range communication between physical neighbours. OFO does not need an accurate grid model nor estimated consumption of non-controllable loads, affords fast calculations, and demonstrates robustness to uncertainties and disturbances, which render it particularly suitable for real-time distribution system applications. However, many OFO controllers require centralized communication, making them susceptible to single-point failures. This paper proposes a distributed OFO design based on a nested feedback optimization strategy and analyzes its convergence. Numerical study results demonstrate that the proposed design achieves effective voltage regulation and outperforms other distributed and local approaches.
@@
Distributed communication, online feedback optimization, reactive power control, voltage regulation
@@ I Introduction
The rapid growth of distributed energy resources (DERs) has introduced significant uncertainty and volatility to distribution systems. With the high resistance/reactance ratios of distribution cables, voltage fluctuations are occurring more frequently. Consequently, real-time voltage regulation has become a significant operational challenge for distribution system operators (DSOs). The conventional approach of DSOs is grid reinforcement, but this demands significant investments, a skilled workforce, and considerable time. During this lengthy process, a complementary solution is to exploit the flexibility of network assets and DERs. In this context, online feedback optimization (OFO) has recently emerged as a promising strategy, e.g. in [1, 2, 3, 4, 5, 6].
OFO utilizes measurements as feedback and employs optimization algorithms as feedback controllers to steer physical systems towards their optimal operating points [7], which are defined by the well-established optimal power flow (OPF) problem. Compared to directly solving the OPF problem in a feedforward manner, OFO does not need an accurate grid model nor data of non-controllable loads, affords fast calculations, and is robust against uncertainties and disturbances in distribution systems due to its feedback-based nature [8, 7, 9]. Figure 1 illustrates feedforward and feedback systems.
@@
Figure 1: Block diagrams of feedforward and feedback systems for voltage regulation using reactive power, where $q$ represents system input (reactive power), $v$ represents system output (voltages), $\omega$ ($\widehat{\omega}$) represents (estimated) system disturbance (non-controllable loads), and $h$ maps system input and disturbance to system output. This figure is adapted from [10].
OFO also brings distinct advantages over other existing approaches. Compared to local droop control [11, 3, 12], OFO ensures more reliable voltage regulation [3] and can pursue grid-level objectives. In contrast to deep reinforcement learning [13, 14], OFO does not need a complicated (often centralized offline) training process while offering theoretical guarantees to obtain globally optimal solutions. Furthermore, unlike using traditional voltage regulators such as on-load tap changers and capacitor banks [15], OFO provides rapid voltage regulation with inverter-interfaced DERs and can complement these slow-reacting devices in real time [16].
While many OFO controllers rely on centralized communication [1, 17, 2, 8, 4, 5, 18, 9, 19, 20, 21], distributed OFO based on communication between physical neighbours appears to be a compelling approach due to its robustness to single-point failures[6, 3, 16, 22, 23, 24]. Among the existing distributed designs, [6, 3, 16] demand particular forms of objective functions, while [22] relies on disseminating global information through neighbouring nodes. However, the systematic delays inherent in this information propagation may raise stability issues, potentially undermining its efficacy in practical distribution systems. In [23], the distributed design was achieved by only keeping components in the network sensitivity matrix related to local and neighbouring nodes. This heuristic approach does not assure adequate voltage regulation.
In this paper, we propose a distributed OFO design inspired by [24], where the distributed implementation was enabled by the sparsity of the inverse of the network sensitivity matrix $\mathbf{X}$, i.e. $\mathbf{X}^{-1}$ is sparse. Both our proposed and their approaches continue with scaling the gradient by $\mathbf{X}^{-1}$ followed by a projection step to ensure satisfaction of the DER capacity constraint. It is conceptually attractive to use the Euclidean distance in the projection, which can be implemented using only local information. However, as shown in [25], the distance measured by the norm $\lVert\mathbf{z}\rVert_{\mathbf{X}}=\sqrt{\mathbf{z}^{\intercal}\mathbf{X}\mathbf{z}}$ should be used instead in the projection step to ensure descent iterations and thus convergence. With the Euclidean distance, it would be a two-metric approach which can lead to algorithm divergence [25]. Solving this non-Euclidean projection problem would however require global information since $\mathbf{X}$ is dense, contradicting the distributed OFO design. As a remedy, [24] introduced additional dualization of the DER capacity constraint on top of the Euclidean projection to ensure the DER constraint satisfaction. Nevertheless, this strategy takes a prohibitively large number of iterations to converge, which may jeopardize its performance in online implementation.
@@
We tackle this challenge by proposing an iterative algorithm to solve this non-Euclidean projection problem that requires only local information and operates in an online feedback-based manner. This leads to our overall nested feedback optimization approach, where the outer loop comprises the OFO iterations requiring communication between physical neighbours and yielding tentative but not necessarily feasible DER setpoints, while the inner loop solves the non-Euclidean projection problem to map the tentative setpoints to actual feasible setpoints using local information. To summarize, the main contributions of this paper are:
  • A local iterative algorithm is proposed to solve the non-Euclidean projection problem that operates in an online feedback-based manner.
  • An efficient nested feedback optimization strategy that ensures adequate voltage regulation while eliminating the need for an accurate grid model and consumption of non-controllable loads is developed.
  • The proposed strategy is theoretically analyzed and numerically assessed on a 96-bus system with high-resolution generation and load data.
@@
The remainder of this paper is structured as follows: Section II presents the system modeling, problem formulation, and a centralized OFO strategy. Section III introduces the proposed nested approach and analyzes its convergence. Section IV presents a numerical study, while Section V draws conclusions and discusses future work.
II Problem formulation and centralized OFO
II-A Distribution system modeling
[…]
The linearized DistFlow equations in (1) were proposed in [26] and can be used to model balanced radial distribution systems. Equations (1a)-(1b) represent nodal active and reactive power balance constraints, respectively. The voltage relation is modeled in (1c), which further leverages the assumption that $v_{i}^{2}-v_{j}^{2}\approx 2(v_{i}-v_{j})$ since $v_{i}\approx 1\text{ pu},\forall i\in\mathcal{N}$.
$P_{ij}+p_{j}=\sum_{k:(j,k)\in\mathcal{E}}P_{jk}+p^{d}_{j},\forall j\in\mathcal{N}^{+},$ (1a)
$Q_{ij}+q_{j}=\sum_{k:(j,k)\in\mathcal{E}}Q_{jk}+q^{d}_{j},\forall j\in\mathcal{N}^{+},$ (1b)
$v_{i}-v_{j}=r_{ij}P_{ij}+x_{ij}Q_{ij},\forall(i,j)\in\mathcal{E}.$ (1c)
Applying (1) directly in a feedforward voltage regulation scheme requires estimates of $p_{j}^{d}$ and $q^{d}_{j}$, $\forall j\in\mathcal{N}^{+}$. Moreover, the modeling inaccuracy will also result in voltage approximation errors. To synthesize the proposed feedback controller for voltage regulation, we leverage (1) to derive network sensitivities, which relate changes in nodal active and reactive power injections to voltage variations. Following the steps in [27] or [28], the linear relation in (2) can be constructed.
$\mathbf{v}=v_{0}\mathbf{1}+\mathbf{R}(\mathbf{p}-\mathbf{p}^{d})+\mathbf{X}(\mathbf{q}-\mathbf{q}^{d}).$ (2)
The $N\times N$-dimensional symmetrical matrices $\mathbf{R}$ and $\mathbf{X}$ are given in (3), where $\mathcal{E}_{i}$ represents the set of cables on the unique path from the substation bus to bus $i$.
$\mathbf{R}_{ij}=\sum_{(h,k)\in\mathcal{E}_{i}\cap\mathcal{E}_{j}}r_{hk},\mathbf{X}_{ij}=\sum_{(h,k)\in\mathcal{E}_{i}\cap\mathcal{E}_{j}}x_{hk}.$ (3)
These matrices accordingly capture the network sensitivities, i.e.
$\frac{\partial\mathbf{v}}{\partial\mathbf{p}}=\mathbf{R},\frac{\partial\mathbf{v}}{\partial\mathbf{q}}=\mathbf{X}.$ (4)
When the resistances and reactances of the cables are all positive, [27] shows that $\mathbf{R}$ and $\mathbf{X}$ are both symmetrical positive definite. Furthermore, [24] shows that their inverse matrices $\mathbf{R}^{-1}$ and $\mathbf{X}^{-1}$ have the sparsity pattern such that
$\mathbf{R}_{ij}^{-1}\neq 0,\mathbf{X}_{ij}^{-1}\neq 0\iff(i,j)\in\mathcal{E}\text{ or }i=j.$ (5)
Figure 2 visualizes $\mathbf{X}$ and $\mathbf{X}^{-1}$ for the test system in Section IV, where the sparsity patterns are shown.
(a) $\mathbf{X}$ (pu/kVar)
(b) $\mathbf{X}^{-1}$ (kVar/pu)
Figure 2: Heat maps of the network sensitivity matrix $\mathbf{X}$ and its inverse $\mathbf{X}^{-1}$.
II-B Problem formulation
@@
Consider the problem of distribution system voltage regulation using reactive power as formulated in (6). The objective function (6a) minimizes the reactive power use cost, where $c_{i}>0$ is the cost parameter. Constraint (6b) enforces the lower and upper voltage limits $\underline{v}$ and $\overline{v}$, where $\mu_{i}$ and $\lambda_{i}$ are the respective dual variables. Finally, (6c) ensures that the reactive power injection $q_{i}$ respects the lower and upper limits $\underline{q}_{i}$ and $\overline{q}_{i}$, which capture DER capacity and power factor constraints. It is noteworthy that $v_{i}$ is dependent on the decision variable $q_{i}$ and other disturbances in the system such as active and reactive power consumption of non-controllable loads.
$\underset{q_{i},\forall i\in\mathcal{N}^{+}}{\operatorname{minimize}}$ $\quad\sum_{i\in\mathcal{N}^{+}}\frac{1}{2}c_{i}q_{i}^{2},$ (6a)
$\operatorname{s.t.}$ $\quad\underline{v}\leq v_{i}\leq\overline{v}:\mu_{i},\lambda_{i},\forall i\in\mathcal{N}^{+},$ (6b)
$\underline{q}_{i}\leq q_{i}\leq\overline{q}_{i},\forall i\in\mathcal{N}^{+}.$ (6c)
For further developments, the dual problem of (6) is formulated in (7).
@@
$\underset{\mu_{i}\geq 0,\lambda_{i}\geq 0,\forall i\in\mathcal{N}^{+}}{\operatorname{maximize}}\big{\{}\underset{\underline{q}_{i}\leq q_{i}\leq\overline{q}_{i},,\forall i\in\mathcal{N}^{+}}{\operatorname{minimize}}\quad\mathcal{L}(\mathbf{q},\boldsymbol{\mu},\boldsymbol{\lambda})\big{\}},$ (7)
where the partial Lagrangian function is defined in (8).
$\mathcal{L}(\mathbf{q},\boldsymbol{\mu},\boldsymbol{\lambda})=\sum_{i\in\mathcal{N}^{+}}\left[\frac{1}{2}c_{i}q_{i}^{2}+\lambda_{i}\left(v_{i}-\overline{v}\right)+\mu_{i}\left(\underline{v}-v_{i}\right)\right].$ (8)
II-C Centralized primal-dual gradient projection algorithm
@@
Solving (6) offline in a feedforward way demands an accurate grid model and disturbance data to evaluate $v_{i}$, which might not be readily available. The feedforward approach also lacks robustness [7]. Thus, this section introduces a feedback controller based on OFO. The controller leverages primal-dual gradient projection (PDGP) as the underlying algorithm [1, 17, 2] to track the time-varying optimizers of (6), which requires a centralized gather-and-broadcast communication architecture. Specifically, the PDGP algorithm performs projected gradient ascent and descent steps iteratively for the dual and primal variables, respectively. At each time step (iteration) $k$, it repeats the following three steps:
@@
    For each bus $i\in\mathcal{N}^{+}$, collect its voltage measurement $\tilde{v}_{i}^{k}$ and update $\lambda_{i}$ and $\mu_{i}$ using (9a)-(9b), where the projection operator is defined as $[u]^{+}=\max(u,0)$, $\alpha$ with various superscripts represents step sizes, while $r^{d}$ and $r^{p}$ in (9c) are the dual and primal regularization factors [1, 17, 2], respectively.
    $\lambda_{i}^{k+1}=\big{[}\lambda_{i}^{k}+\alpha^{d}(\tilde{v}_{i}^{k}-\overline{v}-r^{d}\lambda_{i}^{k})\big{]}^{+},$ (9a)
    $\mu_{i}^{k+1}=\big{[}\mu_{i}^{k}+\alpha^{d}(\underline{v}-\tilde{v}_{i}^{k}-r^{d}\mu_{i}^{k})\big{]}^{+}.$ (9b)
@@
    The DSO broadcasts the updated dual variables $\boldsymbol{\lambda}^{k+1}$ and $\boldsymbol{\mu}^{k+1}$ to all buses.
@@
    For each bus $i\in\mathcal{N}^{+}$, update its reactive power setpoints $q_{i}$ using (9c), where $[u]_{\underline{u}}^{\overline{u}}$ represents the Euclidean projection into $[\underline{u},\overline{u}]$. Since $\mathbf{X}$ is dense, Step 2 is needed to provide information for Step 3.
    $q_{i}^{k+1}=\big{[}q_{i}^{k}-\alpha[c_{i}q_{i}^{k}+[\mathbf{X}(\boldsymbol{\lambda}^{k+1}-\boldsymbol{\mu}^{k+1}+r^{p}\mathbf{q}^{k})]_{i}]\big{]}_{\underline{q}_{i}^{k}}^{\overline{q}_{i}^{k}}.$ (9c)
III Distributed OFO
III-A Two-metric distributed approach
@@
To avoid the centralized communication requirement and enable the distributed communication architecture where communication is only between physical neighbours, [24] proposed to scale the gradient in (9c) by the sparse positive definite matrix $\mathbf{X}^{-1}$, which has non-zero entries only in the diagonal and the $i$-th row $j$-th column if buses $i$ and $j$ are adjacent, i.e. as in (5). Continuing with the Euclidean projection yields (10) which replaces (9c) in the OFO, where $\mathbf{C}$ is a $N$-dimensional diagonal matrix with $c_{i}$ defined in (6a).
$q_{i}^{k+1}=\big{[}q_{i}^{k}-\alpha[[\mathbf{X}^{-1}\mathbf{C}\mathbf{q}^{k}]_{i}+\lambda_{i}^{k+1}-\mu_{i}^{k+1}+r^{p}q_{i}^{k}]\big{]}_{\underline{q}_{i}^{k}}^{\overline{q}_{i}^{k}}.$ (10)
Computing (10) for node $i$ now requires only information exchange with its physical neighbours. However, the two-metric approach does not necessarily lead to algorithm convergence since it is not in general a descent iteration [25]. This is illustrated in Fig. 3 with an example. For further developments, we define
$\mathbf{\dot{q}}^{k+1}=$ (11)
$\quad\left[q_{i}^{k}-\alpha[[\mathbf{X}^{-1}\mathbf{C}\mathbf{q}^{k}]_{i}+\lambda_{i}^{k+1}-\mu_{i}^{k+1}+r^{p}q_{i}^{k}],i\in\mathcal{N}^{+}\right],$
which are tentative but not necessarily feasible setpoints to be projected for all nodes, and define
$\mathcal{X}^{k}=[\underline{q}_{1}^{k},\overline{q}_{1}^{k}]\times[\underline{q}_{2}^{k},\overline{q}_{2}^{k}]\times\cdots\times[\underline{q}_{N}^{k},\overline{q}_{N}^{k}],$ (12)
where $\times$ is the Cartesian product of sets.
Figure 3: An illustration showing that the two-metric approach is not a descent iteration, where $\mathbf{x}^{k}$ is the current iterate which is also the optimum, $X$ is the feasible region, and $\nabla f^{k}$ is the gradient at $\mathbf{x}^{k}$. With (plain) gradient projection on the left panel, the iterate stays at the optimum, i.e. $\mathbf{x}^{k+1}=\mathbf{x}^{k}$. With the two-metric approach on the right panel where $\mathbf{Y}$ represents the gradient scaling metric, $\mathbf{x}^{k+1}$ leaves the optimum.
III-B Proposed distributed approach